OpenStudy (anonymous):

Solve: −4(x + 10) − 6 = −3(x − 2)

OpenStudy (anonymous):

First Use distributive property

OpenStudy (anonymous):

x= -52

OpenStudy (katrinakaif):

Simplify: -4x - 40 - 6 = -3x + 6

OpenStudy (anonymous):

so it will be : -4x+40-6=-3x+6

OpenStudy (katrinakaif):

Collect like terms: -4x - 46 = -3x + 6 Bring the x's to one side -46 = 1x + 6 -52 = 1x x = -52

OpenStudy (katrinakaif):

A medal would be appreciated. ;)